SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

   
Prev Next


Date: February 25, 2008
Shift Taker: Graham Jackson
-----------
General Features:
(theta,phi)
Warm crystals in hitmap	(52,63), (56,63), (51,64) (and (53,49) in 79937/8&9)		
Warm crystals in 0.01 GeV hitmap (17,43), (30,45), (32,46), (36,64)
Warm crystals in 0.1 GeV hitmap (30,45), (32,46), (36,64)
Warm crystals in 1 GeV hitmap      

Cool crystal  in hitmap          
Cool crystal  in 0.1 GeV hitmap (29,84) 
Cool crystal  in 1 GeV hitmap    		

[1] =  TC Damage:OOS/OOS ~10^-3 in 0-5 all fibers.  (And usually FLinkA/B/C Timeout is high too).
[2] =  Hot hitmap patch in 0.1GeV plot (~30-35,38-45) & a warm one at (~30-35,100-105). With spikes in the corresponding projections.
[3] = Large variation of digi multiplicity at low energy near the start of the run.

0079907: [1]
0079908: Small blob at log(E) ~ 0.5 to 1 GeV around 700ns to right of peak.  Visible on D2Starplus Trigger line only
0079909: ok.
0079910: ok.
0079911: Short run.
0079912: [1] & 4-10 all fibers.
0079913: ok.
0079914: ok.
0079915: [1]
0079916: ok.
0079917: Short run.
0079918: [1]
0079919: Evidence of noisy crystals spattered about the didi timing peak.  TC Dam OOS ~10^-3 on 1-3 all fibers.
0079926: [2]
0079927: ok.
0079928: [1] [3]
0079929: [2] Short run though.
0079930: ok.
0079931: [2] [3]
0079932: [1]
0079933: [1] & 4-10 all fibers.
0079934: [1]
0079935: ok.
0079936: [1]
0079937: ok.
0079938: ok.
0079939: ok.
0079940: ok.
0079941: ok.
0079942: Evidence of noisy channels on digi timing plot at logE~0 and 1.  TCDamage - OOS on 1-3 all fibers ~10^-2.  V. Short run.
0079943: [1]


Crystal Spectra:
"Badness 1" at theta=37-40, phi=43

Lightpulser:
16 channels dead, 0 channels out of time. Updated on: 25-February 2008 09:05 PST


 

  Links

  • Shift Report Index
  • Back to the Shifter instructions